OpenStudy (anonymous):

What is the rate on an investment that triples $3654 in 28 years? Assume interest is compounded monthly

OpenStudy (ranga):

Here is the compound interest formula: \[\Large A = P(1 + \frac{ r }{ n })^{nt}\]A = Amount at maturity = $3654 * 3 = $10,962 P = Principal Amount = $3654 r = Annual interest rate in decimal = ? n = compounding period (compounded how many times a year) = 12 t = years invested = 28 years. Solve for r. It will be in decimal. Multiply by 100 to convert it into percentage.